# Electromechanical computational model of the human stomach

### v1

@@ 2.2.4 Prestress
From Eq. 15, one can see immediately that the effects of growth and remodeling prior to time $t=0$ that are captured by $\boldsymbol{F}_{\textup{gr}}^{i}$ are exactly the inverse of the elastic prestretch $\boldsymbol{F}_{\textup{e}}^{i}(0)$ in the initial configuration. So, determining $\boldsymbol{F}_{\textup{gr}}^{i}$ is basically identical to determining the elastic prestretch or prestress in the initial configuration. To this end, one has to identify $\boldsymbol{F}_{\textup{e}}^{i}$ such it ensures mechanical equilibrium for the given loading in the initial configuration. Such inverse problems are in general ill-posed, especially if the tissue is modeled as a mixture of several constituents whose elastic deformation in the initial configuration can vary independently (as is the case here because in principle also the $\boldsymbol{F}_{\textup{gr}}^{i}$ can vary independently). One typically needs some physiologically reasonable assumptions to overcome the ill-posedness of such problems. To this end, we rely on the concept of tensional homeostasis. It is well-known that a continuous turnover of mass in collagenous tissues ensures that collagen fibers in healthy tissues typically reside in some mechanical target state, the homeostatic state. Although the precise definition of this state remains under debate, many models assume that its hallmark is a specific homeostatic fiber stress [41]. Herein, we adopt this assumption. That is, we assume that for the fibers in our material, the inelastic deformation gradient $\boldsymbol{F}_{\textup{gr}}^{i}$ is such that the fibers exhibit a specific homeostatic fiber stress. As this fiber stress is easy to translate in a fiber stretch, which again is easy to translate into the respective $\boldsymbol{F}_{\textup{gr}}^{i}$, both $\boldsymbol{F}_{\textup{e}}^{i}$ and $\boldsymbol{F}_{\textup{gr}}^{i}$ are directly determined by the hypothesis that the collagen and muscle fibers in the stomach are in the initial configuration in their homeostatic state. This reduces the inverse problem to determining the $\boldsymbol{F}_{\textup{e}}^{\textup{gm}}$ and $\boldsymbol{F}_{\textup{gr}}^{\textup{gm}}$ of the ground matrix. The ground matrix consists to a large extent of elastin. Elastin forms a highly elastic largely isotropic protein network that enables reversible shape changes. In particular the high water content makes the ground matrix nearly incompressible under the conditions relevant herein. Hence, we model the ground matrix as an isotropic, elastic, incompressible material (cf. A). This directly leads to the assumption that $\boldsymbol{F}_{\textup{gr}}^{\textup{gm}}$ is isochoric and rotation-free, i.e., symmetric.
@@
To solve then the inverse problem, we use the iterative prestress algorithm that is illustrated in Figure 3 and based on [37, 38, 39]. We initialize the algorithm with $\boldsymbol{F}_{\textup{gr}}^{\textup{gm}}=\boldsymbol{I}$. During the first $k_{p}$ iterations, the intraluminal pressure $p$ and fiber prestretches $\lambda_{\textup{e}}^{i},\,i\in{\textup{c},\textup{l}}$ are gradually increased to the known intragastric pressure in vivo and the homeostatic stretch of collagenous fibers. For all subsequent iterations ($k\geq k_{p}$), both are held constant. In each iteration, we use the current guess $\boldsymbol{F}_{\textup{gr},k}^{\textup{gm}}$} to solve the nonlinear equilibrium equations through a Newton-Raphson method. Subsequently, we evaluate the deformation gradient $\boldsymbol{F}_{k}$ and its isochoric part $\overline{\boldsymbol{F}}_{k}=\det(\boldsymbol{F}_{k})^{-\nicefrac{{1}}{{3}}}\boldsymbol{F}_{k}$. Ultimately, we seek to determine $\boldsymbol{F}_{\textup{gr},k}^{\textup{gm}}$} such that the deformation gradient is equal to the identity tensor so that the shape of the reference and initial configuration are identical as required by definition. If $\overline{\boldsymbol{F}}_{k}$ is not yet equal to the identity tensor within a sufficiently small tolerance, we have to modify $\boldsymbol{F}_{\textup{gr}}^{\textup{gm}}$ such that we move closer towards this state in the next iteration step. To achieve this, we have to make sure that the stretches implied by $\boldsymbol{F}_{\textup{gr},k+1}^{\textup{gm}}$} become smaller than the ones implied by $\boldsymbol{F}_{\textup{gr},k}^{\textup{gm}}$} in exactly those principal directions where $\boldsymbol{F}_{k}$ has eigenvalues larger than one and vice versa. This can be achieved by multiplying $\boldsymbol{F}_{\textup{gr}}^{\textup{gm}}$ with the inverse of $\overline{\boldsymbol{F}}_{k}$. Therefore, we compute
@@ 2.2.5 Active strain
Active deformation of the gastric wall arises from the contraction of SMCs. These are directional contractile units that contract only along their fiber direction. In the stomach wall, two primary muscle fiber directions are present, the circumferential and longitudinal direction [5], represented by the unit vectors $\boldsymbol{f}_{\textup{R}}^{\textup{c}}$ and $\boldsymbol{f}_{\textup{R}}^{\textup{l}}$ in the reference configuration. For these two fiber families, we adopt a transversely isotropic, isochoric active deformation gradient
@@
In accordance with [16], we assume that the muscle activation $\gamma$ is a direct function of the SMCs transmembrane voltage $v^{\textup{smc}}$, i.e., $\gamma=\gamma(v^{\textup{smc}})$. While this represents a simplification of the complex excitation–contraction coupling mechanism in gastric SMCs, [16] demonstrated that it yields a sufficiently accurate model of electromechanical coupling for organ-scale simulations. The activation function is given by
@@ 2.3 Electrophysiological model
We employ a phenomenological approach to model gastric electrophysiology [42, 43]. The normalized transmembrane potentials of ICCs and SMCs are represented by the spatio-temporal scalar functions $v^{\textup{icc}}:\Omega_{0}\times[0,\infty)\rightarrow[0,1]$ and $v^{\textup{smc}}:\Omega_{0}\times[0,\infty)\rightarrow[0,1]$, respectively. The propagation of these potentials across the gastric tissue is governed by two coupled monodomain formulations
@@
where $\sigma^{\textup{icc}}$ and $\sigma^{\textup{smc}}$ denote the isotropic diffusion coefficients associated with ICCs and SMCs, respectively. The term $D_{\textup{gap}}$ models the homogenized coupling resistance between the two cell types via gap junctions, and $I_{\textup{ion}}^{\textup{icc}}$ and $I_{\textup{ion}}^{\textup{smc}}$ denote the total transmembrane ionic currents in ICCs and SMCs, respectively. The coefficients $\chi^{\textup{icc}}$, $\chi^{\textup{smc}}$, and $\chi^{\textup{gap}}$ act as dimensionless scaling factors that modulate the relative contributions of transmembrane ($\chi^{\textup{icc}}$, $\chi^{\textup{smc}}$) and intercellular coupling terms ($\chi^{\textup{gap}}$) in the tissue-scale model. In biophysically motivated monodomain formulations, these coefficients represent the surface-to-volume ratio, which naturally emerges from analyzing how membrane surface area contributes to transmembrane ionic and capacitive currents per unit tissue volume, bridging microscopic cellular physiology and macroscopic electrical wave propagation. In contrast, the present phenomenological formulation employs normalized transmembrane potentials and effective ionic current terms. Thus, the coefficients $\chi^{\textup{icc}}$ and $\chi^{\textup{smc}}$ are interpreted as spatially varying weights that reflect regional heterogeneity in cell density. This flexible representation preserves the fundamental ionic kinetics as well-known from cardiac electrophysiology [42]. For the definition of $\chi^{\textup{icc}}$, $\chi^{\textup{smc}}$, and $\chi^{\textup{gap}}$, refer to Section 2.4.3. The formulations of the ionic currents are taken from [16] and are recalled in B.
@@ 2.4 Spatially non-uniform parametrization
The stomach exhibits highly spatially varying features such as fiber orientations [29, 5, 3], intrinsic frequencies [44, 45, 26], and passive material properties [3, 29, 4]. To accommodate this heterogeneity in computational models, particularly in patient-specific geometries, we move from a purely anatomical description based on the fundus, corpus, and antrum to a functionally driven representation of the stomach [5]. Functional regions, such as the proximal and distal stomach, are strongly associated with physiological roles in motility and electrical activity, though they do not map directly onto anatomical landmarks (cf. Figures 1 and 5). In practice, patient-specific geometries often lack detailed microstructural data [29], which limits direct spatial assignment of physiological parameters. To address this, we propose a framework that builds functionally meaningful coordinate systems derived from harmonic scalar fields. These fields are generated by solving Laplace-Dirichlet problems [46] over the reference configuration $\Omega_{\textup{R}}$ and clearly defined anatomical landmarks. They serve as normalized control coordinates for assigning spatially varying model parameters in a way that reflects the underlying functional architecture. This approach enables consistent parameterization across realistic stomach geometries, a critical feature given the substantial inter-subject variability in anatomical shape. Similar approaches have been successfully applied in cardiac modeling to define fiber orientations [47, 48, 49] and principle conduction axes [50]. Here, we extend this method to address a broader range of heterogeneities in the stomach, including, but not limited to, fiber directions in gastric tissue. To this end, we define three harmonic scalar fields as smooth functions of the material point $\boldsymbol{X}\in\Omega_{\textup{R}}$, each aligned with a distinct physiological axis: (i) $\phi_{\textup{ep}}(\boldsymbol{X})$, from the esophagus to the pyloric sphincter, representing the longitudinal direction; (ii) $\phi_{\textup{fp}}(\boldsymbol{X})$, spanning the fundus apex to the pyloric sphincter; and (iii) $\phi_{\textup{gl}}(\boldsymbol{X})$, from the greater to the lesser curvature, aligned with the circumferential direction.
@@ 2.4.1 Principal fiber directions
Peristaltic contractions are chiefly driven by the longitudinal and circumferential muscle fibers. The longitudinal layer extends from the esophagus to the pyloric sphincter, aligning with esophageal muscles and crossing the fundus apex. In contrast, the circumferential fibers run perpendicular to the axis from the fundus to the duodenum [5].
@@
where $\boldsymbol{f}_{\textup{R}}^{\textup{l}}$ and $\boldsymbol{f}_{\textup{R}}^{\textup{c}}$ are the longitudinal and circumferential fiber directions, respectively, and $\boldsymbol{n}$ corresponds to the outward normal direction. The scalar fields and their associated Dirichlet boundary conditions used to construct the fiber directions are presented in Figures 4a and 4b. The computed fiber directions are visualized in Figures 4d and 4e, whereas the outward normal vector is shown in Figure 4f.
@@ 2.4.2 Intrinsic frequencies of ICCs and passive behavior of SMCs
Anatomically, the excitability distribution follows a distinct spatial pattern: ICCs located in the fundus (proximal stomach, above the pacemaker region) exhibit nearly no excitability and thus do not actively contribute to slow-wave initiation. The pacemaker region itself exhibits the highest intrinsic frequency and excitability. Moving away from the pacemaker, excitability and intrinsic frequency decrease gradually both along the longitudinal axis toward the pylorus and circumferentially from the lesser to the greater curvature. In the present model, this spatial variation is captured by an excitability parameter $a^{\textup{icc}}(\boldsymbol{X})$, which controls the local responsiveness of ICCs. We build on our previous work [16], where a similar formulation was introduced for idealized geometries, and extend it here to a realistic stomach geometry. This excitability parameter depends on two scalar fields that define an intrinsic coordinate system aligned with physiologically meaningful axes of the stomach: $\xi_{l}(\boldsymbol{X})\in[0,1]$, representing the longitudinal axis of the stomach, and $\xi_{c}(\boldsymbol{X})\in[0,1]$, representing the circumferential coordinate from the greater to the lesser curvature. Both normalized control coordinates are constructed by solving independent Laplace-Dirichlet problems over the reference configuration $\Omega_{\textup{R}}$. The longitudinal harmonic field $\phi_{\textup{fp}}$ is defined in Eq. 26, while the circumferential field $\phi_{\textup{gl}}$ is obtained by solving
@@
To support spatially varying excitability with physiologically meaningful regional distinctions, we define a threshold value $\phi^{*}\in(0,1)$, corresponding to an isocontour value of $\phi_{\textup{fp}}$. This value is automatically determined from anatomical features, such as points near the esophageal sphincter on the lesser curvature. Using this threshold, we define the normalized longitudinal coordinates $\xi_{l}$ in a piecewise manner as
@@
These coordinates allow spatial modulation of parameters in both regions independently, while preserving anatomical and physiological relevance. Accordingly, the excitability of ICCs is modeled as
@@
where $r\in\{\textup{prox},\textup{distal}\}$ indicates whether the current point $\boldsymbol{X}$ is in the proximal or distal region. The parameter $a_{\textup{min}}^{\textup{icc},r}$ denotes the minimal excitability in each region and $a_{\textup{max}}^{\textup{icc}}$ is the global upper bound for excitability. The shaping functions $f_{l}(\xi_{l})$ and $f_{c}(\xi_{c})$ determine the longitudinal and circumferential variation of excitability in each region:
@@
where the parameters $b_{l}^{r}$, $b_{c}^{r}$, $_{l}^{r}, cc$$_{c}^{r}$ control the shapes and decay behavior in each direction (cf. Table E.1). The resulting excitability field is shown in Figure 4g. This parameter enters the model equations via the inward current terms, where it modulates the ionic fluxes generated by ICCs, see Eq. 49a.
In contrast, SMCs do not actively trigger electrical signals. Therefore, the parameter for SMCs, $a^{\textup{smc}}$, is set to zero ($a^{\textup{smc}}(\boldsymbol{X})=0$), reflecting their passive behavior.

### v2

@@ 2.2.4 Prestress
From Eq. 15, one can see immediately that the effects of growth and remodeling prior to time $t=0$ that are captured by $\boldsymbol{F}_{\textup{gr}}^{i}$ are exactly the inverse of the elastic prestretch $\boldsymbol{F}_{\textup{e}}^{i}(0)$ in the initial configuration. So, determining $\boldsymbol{F}_{\textup{gr}}^{i}$ is basically identical to determining the elastic prestretch or prestress in the initial configuration. To this end, one has to identify $\boldsymbol{F}_{\textup{e}}^{i}$ such it ensures mechanical equilibrium for the given loading in the initial configuration. Such inverse problems are in general ill-posed, especially if the tissue is modeled as a mixture of several constituents whose elastic deformation in the initial configuration can vary independently (as is the case here because in principle also the $\boldsymbol{F}_{\textup{gr}}^{i}$ can vary independently). One typically needs some physiologically reasonable assumptions to overcome the ill-posedness of such problems. To this end, we rely on the concept of tensional homeostasis. It is well-known that a continuous turnover of mass in collagenous tissues ensures that collagen fibers in healthy tissues typically reside in some mechanical target state, the homeostatic state. Although the precise definition of this state remains under debate, many models assume that its hallmark is a specific homeostatic fiber stress [Eichinger2021a]. Herein, we adopt this assumption. That is, we assume that for the fibers in our material, the inelastic deformation gradient $\boldsymbol{F}_{\textup{gr}}^{i}$ is such that the fibers exhibit a specific homeostatic fiber stress. As this fiber stress is easy to translate in a fiber stretch, which again is easy to translate into the respective $\boldsymbol{F}_{\textup{gr}}^{i}$, both $\boldsymbol{F}_{\textup{e}}^{i}$ and $\boldsymbol{F}_{\textup{gr}}^{i}$ are directly determined by the hypothesis that the collagen and muscle fibers in the stomach are in the initial configuration in their homeostatic state. This reduces the inverse problem to determining the $\boldsymbol{F}_{\textup{e}}^{\textup{gm}}$ and $\boldsymbol{F}_{\textup{gr}}^{\textup{gm}}$ of the ground matrix. The ground matrix consists to a large extent of elastin. Elastin forms a highly elastic largely isotropic protein network that enables reversible shape changes. In particular the high water content makes the ground matrix nearly incompressible under the conditions relevant herein. Hence, we model the ground matrix as an isotropic, elastic, incompressible material (cf. A). This directly leads to the assumption that $\boldsymbol{F}_{\textup{gr}}^{\textup{gm}}$ is isochoric and rotation-free, i.e., symmetric.
@@
To solve then the inverse problem, we use the iterative prestress algorithm that is illustrated in Figure 3 and based on [Weisbecker-2014-Prestress, MousaviAvril-2017, Gebauer-2023-CMMCardiac]. We initialize the algorithm with $\boldsymbol{F}_{\textup{gr}}^{\textup{gm}}=\boldsymbol{I}$. During the first $k_{p}$ iterations, the intraluminal pressure $p$ and fiber prestretches $\lambda_{\textup{e}}^{i},\,i\in{\textup{c},\textup{l}}$ are gradually increased to the known intragastric pressure in vivo and the homeostatic stretch of collagenous fibers. For all subsequent iterations ($k\geq k_{p}$), both are held constant. In each iteration, we use the current guess $\boldsymbol{F}_{\textup{gr},k}^{\textup{gm}}$} to solve the nonlinear equilibrium equations through a Newton-Raphson method. Subsequently, we evaluate the deformation gradient $\boldsymbol{F}_{k}$ and its isochoric part $\overline{\boldsymbol{F}}_{k}=\det(\boldsymbol{F}_{k})^{-\nicefrac{{1}}{{3}}}\boldsymbol{F}_{k}$. Ultimately, we seek to determine $\boldsymbol{F}_{\textup{gr},k}^{\textup{gm}}$} such that the deformation gradient is equal to the identity tensor so that the shape of the reference and initial configuration are identical as required by definition. If $\overline{\boldsymbol{F}}_{k}$ is not yet equal to the identity tensor within a sufficiently small tolerance, we have to modify $\boldsymbol{F}_{\textup{gr}}^{\textup{gm}}$ such that we move closer towards this state in the next iteration step. To achieve this, we have to make sure that the stretches implied by $\boldsymbol{F}_{\textup{gr},k+1}^{\textup{gm}}$} become smaller than the ones implied by $\boldsymbol{F}_{\textup{gr},k}^{\textup{gm}}$} in exactly those principal directions where $\boldsymbol{F}_{k}$ has eigenvalues larger than one and vice versa. This can be achieved by multiplying $\boldsymbol{F}_{\textup{gr}}^{\textup{gm}}$ with the inverse of $\overline{\boldsymbol{F}}_{k}$. Therefore, we compute
@@ 2.2.5 Active strain
Active deformation of the gastric wall arises from the contraction of SMCs. These are directional contractile units that contract only along their fiber direction. In the stomach wall, two primary muscle fiber directions are present, the circumferential and longitudinal direction [DiNatale-2023-FunctionalAndAnatomicalRegions], represented by the unit vectors $\boldsymbol{f}_{\textup{R}}^{\textup{c}}$ and $\boldsymbol{f}_{\textup{R}}^{\textup{l}}$ in the reference configuration. For these two fiber families, we adopt a transversely isotropic, isochoric active deformation gradient
@@
In accordance with [brandstaeter2018a], we assume that the muscle activation $\gamma$ is a direct function of the SMCs transmembrane voltage $v^{\textup{smc}}$, i.e., $\gamma=\gamma(v^{\textup{smc}})$. While this represents a simplification of the complex excitation–contraction coupling mechanism in gastric SMCs, [brandstaeter2018a] demonstrated that it yields a sufficiently accurate model of electromechanical coupling for organ-scale simulations. The activation function is given by
@@ 2.3 Electrophysiological model
We employ a phenomenological approach to model gastric electrophysiology [Keener-2009-MathematicalPhysiology, Pullan-2006-ModellingElectricalActivityHeart]. The normalized transmembrane potentials of ICCs and SMCs are represented by the spatio-temporal scalar functions $v^{\textup{icc}}:\Omega_{0}\times[0,\infty)\rightarrow[0,1]$ and $v^{\textup{smc}}:\Omega_{0}\times[0,\infty)\rightarrow[0,1]$, respectively. The propagation of these potentials across the gastric tissue is governed by two coupled monodomain formulations
@@
where $\sigma^{\textup{icc}}$ and $\sigma^{\textup{smc}}$ denote the isotropic diffusion coefficients associated with ICCs and SMCs, respectively, with units in $\mathrm{m}\mathrm{m}^{2}\mathrm{/}\mathrm{s}$. These coefficients are modeled as spatially heterogeneous, isotropic scalar fields, reflecting that electrical conduction is dominated by ICC-mediated propagation in the myenteric plexus, which lacks a preferred fiber orientation [huizinga2009a, Sanders-2016-RegulationSMCFunction], while SMC activation is mainly driven by coupling from ICCs rather than direct fiber-aligned conduction [OGrady-2012-AbnormalInitiationSW, du2013c]. The term $D_{\textup{gap}}$ models the homogenized coupling resistance between the two cell types via gap junctions, and $I_{\textup{ion}}^{\textup{icc}}$ and $I_{\textup{ion}}^{\textup{smc}}$ denote the total transmembrane ionic currents in ICCs and SMCs, respectively. The coefficients $\chi^{\textup{icc}}$, $\chi^{\textup{smc}}$, and $\chi^{\textup{gap}}$ act as dimensionless scaling factors that modulate the relative contributions of transmembrane ($\chi^{\textup{icc}}$, $\chi^{\textup{smc}}$) and intercellular coupling terms ($\chi^{\textup{gap}}$) in the tissue-scale model. In biophysically motivated monodomain formulations, these coefficients represent the surface-to-volume ratio, which naturally emerges from analyzing how membrane surface area contributes to transmembrane ionic and capacitive currents per unit tissue volume, bridging microscopic cellular physiology and macroscopic electrical wave propagation. In contrast, the present phenomenological formulation employs normalized transmembrane potentials and effective ionic current terms. Thus, the coefficients $\chi^{\textup{icc}}$ and $\chi^{\textup{smc}}$ are interpreted as spatially varying weights that reflect regional heterogeneity in cell density. This flexible representation preserves the fundamental ionic kinetics as well-known from cardiac electrophysiology [Keener-2009-MathematicalPhysiology]. For the definition of $\chi^{\textup{icc}}$, $\chi^{\textup{smc}}$, and $\chi^{\textup{gap}}$, refer to Section 2.4.3. The formulations of the ionic currents ($I_{\textup{ion}}^{\textup{icc}}$ and $I_{\textup{ion}}^{\textup{smc}}$) are taken from [brandstaeter2018a] and are recalled in B. In particular, each cell type (ICC and SMC) is modeled using a modified two-variable Mitchell–Schaeffer cell model [MitchellSchaeffer-2003-TwoCurrentModel], which provides a sufficient description of cellular dynamics, while allowing stable and robust long-term simulations of slow-wave propagation, as demonstrated in our previous work [brandstaeter2018a].
@@ 2.4 Spatially non-uniform parametrization
The stomach exhibits highly spatially varying features such as fiber orientations [Patel-Gizzi-2022-GastrointestinalTissueReview, DiNatale-2023-FunctionalAndAnatomicalRegions, Friis-2023-GastricTissue], intrinsic frequencies [vanHelden-2009_GenerationSW, OGrady-2012-AbnormalInitiationSW, EgbujiGrady-2010-OriginPropagationSW], and passive material properties [Friis-2023-GastricTissue, Patel-Gizzi-2022-GastrointestinalTissueReview, Holzer-Stock2025]. To accommodate this heterogeneity in computational models, particularly in patient-specific geometries, we move from a purely anatomical description based on the fundus, corpus, and antrum to a functionally driven representation of the stomach [DiNatale-2023-FunctionalAndAnatomicalRegions]. Functional regions, such as the proximal and distal stomach, are strongly associated with physiological roles in motility and electrical activity, though they do not map directly onto anatomical landmarks (cf. Figures 1 and 6). In practice, patient-specific geometries often lack detailed microstructural data [Patel-Gizzi-2022-GastrointestinalTissueReview], which limits direct spatial assignment of physiological parameters. To address this, we propose a framework that builds functionally meaningful coordinate systems derived from harmonic scalar fields. These fields are generated by solving Laplace-Dirichlet problems [Bayer-2005-LDFiber] over the reference configuration $\Omega_{\textup{R}}$ and clearly defined anatomical landmarks. They serve as normalized control coordinates for assigning spatially varying model parameters in a way that reflects the underlying functional architecture. This approach enables consistent parameterization across realistic stomach geometries, a critical feature given the substantial inter-subject variability in anatomical shape. Similar approaches have been successfully applied in cardiac modeling to define fiber orientations [Piersanti-2021-CardiacFibers, Baier-2020-CardiacFiber, Bayer-2012-RuleBasedFiberHEart] and principle conduction axes [sathar2015a]. Here, we extend this method to address a broader range of heterogeneities in the stomach, including, but not limited to, fiber directions in gastric tissue. To this end, we define three harmonic scalar fields as smooth functions of the material point $\boldsymbol{X}\in\Omega_{\textup{R}}$, each aligned with a distinct physiological axis: (i) $\phi_{\textup{ep}}(\boldsymbol{X})$, from the esophagus to the pyloric sphincter, representing the longitudinal direction; (ii) $\phi_{\textup{fp}}(\boldsymbol{X})$, spanning the fundus apex to the pyloric sphincter; and (iii) $\phi_{\textup{gl}}(\boldsymbol{X})$, from the greater to the lesser curvature, aligned with the circumferential direction.
@@ 2.4.1 Principal fiber directions
Peristaltic contractions are chiefly driven by the longitudinal and circumferential muscle fibers. The longitudinal layer extends from the esophagus to the pyloric sphincter, aligning with esophageal muscles and crossing the fundus apex. In contrast, the circumferential fibers run perpendicular to the axis from the fundus to the duodenum [DiNatale-2023-FunctionalAndAnatomicalRegions].
@@
where $\boldsymbol{f}_{\textup{R}}^{\textup{l}}$ and $\boldsymbol{f}_{\textup{R}}^{\textup{c}}$ are the longitudinal and circumferential fiber directions, respectively, and $\boldsymbol{n}$ corresponds to the outward normal direction. Note that, contrary to engineering intuition, the circumferential and longitudinal directions are not necessarily orthogonal. These directions are defined anatomically, based on experimental observations. To accurately represent them, a second coordinate system is introduced to capture the anatomical circumferential and longitudinal orientations. The scalar fields and their associated Dirichlet boundary conditions used to construct the fiber directions are presented in Figures 4a and 4b. The computed fiber directions are visualized in Figures 4d and 4e, whereas the outward normal vector is shown in Figure 4f.
@@ 2.4.2 Intrinsic frequencies of ICCs and passive behavior of SMCs
Anatomically, the excitability distribution follows a distinct spatial pattern: ICCs located in the fundus (proximal stomach, above the pacemaker region) exhibit nearly no excitability and thus do not actively contribute to slow-wave initiation. The pacemaker region itself exhibits the highest intrinsic frequency and excitability. Moving away from the pacemaker, excitability and intrinsic frequency decrease gradually both along the longitudinal axis toward the pylorus and circumferentially from the lesser to the greater curvature. In the present model, this spatial variation is captured by an excitability parameter $a^{\textup{icc}}(\boldsymbol{X})$, which controls the local responsiveness of ICCs. We build on our previous work [brandstaeter2018a], where a similar formulation was introduced for idealized geometries, and extend it here to a realistic stomach geometry. This excitability parameter depends on two scalar fields that define an intrinsic coordinate system aligned with physiologically meaningful axes of the stomach: $\xi_{l}(\boldsymbol{X})\in[0,1]$, representing the longitudinal axis of the stomach, and $\xi_{c}(\boldsymbol{X})\in[0,1]$, representing the circumferential coordinate from the greater to the lesser curvature. Both normalized control coordinates are constructed by solving independent Laplace-Dirichlet problems over the reference configuration $\Omega_{\textup{R}}$. The longitudinal harmonic field $\phi_{\textup{fp}}$ is defined in Eq. 26, while the circumferential field $\phi_{\textup{gl}}$ is obtained by solving
@@
To support spatially varying excitability with physiologically meaningful regional distinctions, we define a threshold value $\phi^{*}\in(0,1)$, corresponding to an isocontour value of $\phi_{\textup{fp}}$. This value is automatically determined from anatomical features, such as points near the esophageal sphincter on the lesser curvature. Specifically, the geometric center and average radius of the esophageal sphincter outlet are computed. The algorithm then identifies the closest node along the lesser curvature whose radial distance from the center exceeds $1.3$ times the average outlet radius. The value of $\phi_{\textup{fp}}$ at this node is then assigned as the threshold $\phi^{*}$. Consequently, $\phi^{*}$ depends on the specific geometry under consideration; for the geometry in Section 4.2.4, this yields $\phi^{*}\approx 0.543$. Using this threshold, we define the normalized longitudinal coordinates $\xi_{l}$ in a piecewise manner as
@@
Figure 5: Control variables used for spatial parameter modulation visualized from two perspectives. (a) The normalized longitudinal coordinate $\xi_{l}$, defined piecewise in Eq. 29, distinguishes the proximal and distal regions of the stomach and allows longitudinal variation of model parameters. (b) The normalized circumferential coordinate $\xi_{c}$ follows the longitudinal field $\phi_{\textup{gl}}$ and allows circumferential parameter modulation. These coordinates provide a structured framework to spatially vary parameters such as excitability along anatomically meaningful axes.
Although $\xi_{l}$ is piecewise-defined for the proximal and distal stomach, it is constructed to remain continuous across the transition at $\phi_{\textup{fp}}=\phi^{*}$. Both control variables $\xi_{l}$ and $\xi_{c}$ are visualized in Figure 5 and serve to systematically modulate tissue properties in the model along the longitudinal and circumferential directions. These coordinates allow continuous spatial modulation of parameters in both regions independently, while preserving anatomical and physiological relevance. Accordingly, the excitability of ICCs is modeled as
@@
where $r\in\{\textup{prox},\textup{distal}\}$ indicates whether the current point $\boldsymbol{X}$ is in the proximal or distal region. The parameter $a_{\textup{min}}^{\textup{icc},r}$ denotes the minimal excitability in each region and $a_{\textup{max}}^{\textup{icc}}$ is the global upper bound for excitability. The excitability parameter $a^{\textup{icc}}$ enters the model equations via the inward current terms, where it modulates the ionic fluxes generated by ICCs, see Eq. 49a. The shaping functions $f_{l}(\xi_{l})$ and $f_{c}(\xi_{c})$ determine the longitudinal and circumferential variation of excitability in each region:
@@
where the parameters $b_{l}^{r}$, $b_{c}^{r}$, $_{l}^{r}, cc$$_{c}^{r}$ control the shapes and decay behavior in each direction (cf. Table F.1). The resulting excitability field is shown in Figure 4g. In contrast, SMCs do not actively trigger electrical signals. Therefore, the parameter for SMCs, $a^{\textup{smc}}$, is set to zero ($a^{\textup{smc}}(\boldsymbol{X})=0$), reflecting their passive behavior.